OpenStudy (anonymous):

Write and solve an equation to find the answer for the problem below. In Mrs. Narain's classroom, students sit at tables and desks. A table seats 4 students. A desk seats 1 student. There are twenty-five students in the class. If 9 students sit at desks, how many tables are used? 4 8 sixteen twenty-one

OpenStudy (nincompoop):

25 = 4t + 9d look for t, where d = 1

OpenStudy (anonymous):

thanks

OpenStudy (nincompoop):

25 = 4t + 9(1) 25 = 4t + 9 subtract 9 on both sides to leave out 4t. since we want t we isolate it. 25 - 9 = 4t + 9 - 9 16 = 4t divide by 4 to get t

OpenStudy (anonymous):

kk

OpenStudy (nincompoop):

i just woke up. going to make breakfast ttyl

OpenStudy (anonymous):

kk bye lol